a) Determine the binding energy per nucleon of a nucleus given the following data

mass C12 nucleus = 11.99670u, mass of a proton=1.00728u, mass of neutron=1.00867. 1u=931Mev

 

b) Sketch the curve for binding energy per nucleon against mass number.

Relate this to the energy released in nuclear fusion and fission


1
Expert's answer
2021-02-15T17:40:44-0500

a)


"BE=\frac{1}{12}(6\cdot1.00728+6\cdot1.00867-11.99670)931\\BE=7.68\frac{MeV}{nucleon}"

b)